# Efficiently Training 7B LLM with 1 Million Sequence Length on 8 GPUs

### v1

@@ 4.2. Bi-level Memory Planning
The offline DSA problem assumes a predefined sequence of memory allocations and deallocations, and aims to determine the address of each allocated memory block and at the same time minimizing the peak memory usage. Parameters of offline DSA problem includes:
@@
Here the first constraint and the last two constraints define and limit peak memory, while the second and third constraints ensure non-overlapping tensors. Following this formulation, the solution for each tensor’s address is optimal. However, modern LLM training involves thousands of allocation and deallocation requests, which makes this MIP problem computationally intractable.
Fortunately, all transformer layers have identical structures and memory request sequences, which presents repetitive substructures within the MIP problem. By leveraging this inherent repetitiveness, we can devise a bi-level hierarchical MIP optimization algorithm, which is both computationally feasible and effective.
@@
As shown in the bottom of Figure 8, we first solve the offline DSA sub-problem for just one transformer layer’s forward (backward) pass, which is called the first-level MIP. This offline DSA problem can be simply solved by any MIP solver. After this step, the peak memory needed for the forward (backward) propagation of a single transformer layer, as well as the address of each transient tensor within a transformer layer is determined. After solving the sub-problem for one transformer layer, all other transformer layers can reuse the same memory address for (de)allocation.
Subsequently, we can replace the original fine-grained memory request sequence of a transformer layer forward (backward) propagation with a “pseudo” large memory request pair, as shown in Figure 8. After the substitution, this reformulated memory request sequence also satisfies the formulation of an offline DSA problem, with a size small enough to be efficiently solved. We then leverage the MIP solver again to solve this second-level MIP problem. After this step, the addresses of all activation tensors, as well as the peak memory needed for all transient activation tensors can be determined.
@@ 4.3.1. Overview
Figure 10 illustrates the overall architecture of Memo. First, the job profiler takes in the model configuration and user-defined settings, then executes a training iteration to profile the memory requests directed to the PyTorch CUDA allocator during the training phase. The job profiler also determines offloading fraction $\alpha$ by solving optimization problem in Section 4.1. These memory requests comprise a sequence of allocation and deallocation instructions. Afterwards, the memory planner receives the memory requests, executes the bi-level MIP optimization algorithm and, generates a memory plan, which constitutes the addresses of all activation tensors during one training iteration. Finally, the runtime executor reads the memory plan and conducts the training process.
@@ 4.3.2. Job Profiler
However, naively recording all memory requests may lead to OOM error. For example, profiling a GPT-7B model with a sequence length of 512K on 8 GPUs can result in OOM error since the proposed techniques are not employed in profiling. Fortunately, as the transformer layers have identical memory footprint, we leverage this property by only profiling a single transformer layer’s memory footprint and then applying it to all transformer layers.
When the sequence is too long, we cannot even profile one single transformer layer. In such extreme cases, we turn to the CUDA Unified Memory feature, which enables the swapping between GPU memory and CPU memory under the hook, effectively creating an illusion of unlimited GPU memory. By integrating CUDA Unified Memory support into the PyTorch CUDA allocator, we have successfully managed to profile the training of extremely long context lengths without encountering OOM errors.
The profiler also gathers the basic information to determine $\alpha$ in Section 4.1, including the size of each skeletal activation tensor, and the forward time of a single transformer layer. Subsequently, it solves for the optimal $\alpha$ to maximize the overlapping of computation and tranmission as well as to avoid CPU OOM error.
@@ 4.3.3. Memory Planner
Given the memory request sequence generated by the job profiler, memory planner executes the bi-level MIP optimization algorithm as introduced in Section 4.2 to generate a memory plan, which includes the address of each activation tensor and the peak memory usage needed during training. In all our experiments, memory planning takes less than 5 minutes, which is negligible compared to the training time of LLMs.
@@ 4.3.4. Runtime Executor
The runtime executor takes the memory plan as input, and executes the training process. It is built on the top of Megatron-LM (Shoeybi et al., 2019) and TransformerEngine (NVIDIA, 2024b), one of the most popular LLM training frameworks. The runtime executor utilizes two rounding buffers for the storage of skeletal activations, as introduced in Section 4.1. Meanwhile, the transient activation tensors are allocated and discarded according to the memory plan.
Three CUDA streams are employed for efficient overlapping of data transmission and GPU computation, which are for GPU computation, activation offloading from GPU to CPU, and activation prefetching from CPU to GPU, respectively. Figure 11 shows the scheduling of computation and transmission. After the computation of a transformer layer’s forward pass, the skeletal activations of this layer are scheduled to be transferred to the CPU memory, which can overlap with the computation of the next layer. Before the backward computation of a transformer layer, the forward skeletal activations of the previous layer are scheduled to be fetched back to GPU. In addition, token-wise tensor recomputation is also scheduled before the layer’s backward pass. By hiding the activation swapping with computation and enabling the fractional, token-wise activation recomputation, Memo minimizes the overhead of activation rematerialization at full stretch.
5. Experiments
In this section, we conduct experiments across various model sizes and input sequence lengths to show that Memo achieves superior efficiency in longer context training of LLMs.
5.1. Setup
Baselines: We select two widely-used LLM training frameworks as baselines for our experiments. The first is Megatron-LM (commit id: ccfeda47cb) (Shoeybi et al., 2019) in conjunction with TransformerEngine (v1.3) (NVIDIA, 2024b). Megatron-LM, maintained by NVIDIA, is renowned for its comprehensive support of hybrid parallelisms, including DP, TP, PP, SP, and CP. The other baseline is Megatron-Deepspeed (commit id 7eb36a11b3) paired with DeepSpeed (v0.14.3) (Rasley et al., 2020), which is recognized for ZeRO optimizers and DeepSpeed-Ulysses (Jacobs et al., 2023), a novel parallel training strategy designed for long context LLM training.
Metrics: We use two important evaluation metrics to measure the training efficiency, which are Model FLOPs Utilization (MFU) and Tokens per GPU per Second (TGS). MFU is defined as the ratio of model FLOPs per second to the theoretical peak FLOPs per second of the GPU (e.g. 312 TFLOPS for NVIDIA A800 GPUs) (Korthikanti et al., 2022). Based on FlashAttention (Dao, 2023) and considering the causal mask, the exact formula for calculating model FLOPS per sample is:
$6\cdot s\cdot P+6\cdot n\cdot h\cdot s^{2}.$
MFU is a standard metric that measures the training efficiency of how model FLOPs utilize computational resources. On the other hand, TGS directly measures training throughput, providing a clear view of how quickly a model can be trained using a given amount of training samples. Both metrics are crucial for LLM researchers and engineers, enabling comparisons among various training strategies (including distributed parallelisms and activation recomputation).
@@ 5.1. Setup
Workloads: Our experiments cover a wide range of workloads to examine the strength of Memo. In particular, we consider training the 7B, 13B, 30B and 65B GPT models on 8, 16, 32, and 64 GPUs, respectively, with various sequence lengths ranging from 64K to 1408K. The detailed model configurations are shown in Table 2.
@@
Table 3. MFU (Model FLOPS Utilization) and TGS (Throughput per GPU per Second) of different methods on different number of GPUs, model sizes, and sequence lengths. “DS” stands for “DeepSpeed”, and “Mega” stands for “Megatron-LM”. ✗oom stands for Out Of Memory error when the system runs out of GPU memory, and ✗oohm stands for Out Of Host Memory error, which means the CPU memory is depleted.
Sequence Length
# GPUs Model Size Method 64K 128K 256K 384K 512K 640K 768K 896K 1024K 1152K 1280K 1408K
27.95% 25.46% 23.38% oom oom oom oom oom oom oom oom oom
DS 935.76 555.51 296.48 oom oom oom oom oom oom oom oom oom
41.55% 24.13% 29.07% 27.98% 34.43% 30.90% oom oom oom oom oom oom
8 7B Mega 1417.74 526.62 368.65 250.07 237.56 173.63 oom oom oom oom oom oom
52.34% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71% 52.30% oom oom oom
Memo 1786.22 1111.99 679.92 474.02 357.70 295.51 245.76 215.98 188.73 oom oom oom
27.97% 25.45% 21.98% oom oom oom oom oom oom oom oom oom
DS 553.85 333.51 171.82 oom oom oom oom oom oom oom oom oom
38.51% 23.02% 25.30% 22.88% 29.10% 19.41% oom oom oom oom oom oom
16 13B Mega 762.46 301.67 197.78 127.40 125.86 83.94 oom oom oom oom oom oom
52.65% 50.93% 51.22% 51.91% 52.40% 52.13% 51.71% 51.76% 52.06% 51.74% 51.78% 52.10%
Memo 1042.5 667.41 400.39 289.11 226.65 184.33 154.63 134.08 118.96 105.74 95.72 87.93
29.93% 25.54% oom oom oom oom oom oom oom oom oom oom
DS 296.41 176.92 oom oom oom oom oom oom oom oom oom oom
35.76% 14.70% 17.15% 23.32% oom oom oom oom oom oom oom oom
32 30B Mega 354.20 101.85 74.24 73.37 oom oom oom oom oom oom oom oom
52.12% 49.66% 50.00% 50.69% 51.06% 51.72% 51.18% 51.50% 51.24% 51.73% 51.59% oohm
Memo 516.16 344.09 216.41 159.52 126.23 105.28 88.55 77.48 68.20 61.72 55.79 oohm
31.05% 26.13% 22.07% 20.40% 19.83% 19.06% 19.53% 19.12% 19.00% 19.11% 18.90% oom
DS 149.8 90.15 48.51 32.89 25.24 20.04 17.50 14.93 13.14 11.86 10.64 oom
22.79% 15.10% 9.57% 12.07% 5.32% oom oom oom oom oom oom oom
64 65B Mega 109.94 52.10 21.03 19.46 6.77 oom oom oom oom oom oom oom
47.80% 48.61% 49.87% 48.85% 49.71% 50.05% 51.16% 51.05% 51.27% 51.20% 51.42% 51.45%
Memo 230.62 167.69 109.58 78.76 63.29 52.65 45.84 39.85 35.45 31.77 28.94 26.50
@@ 5.2. End-to-end Evaluation
We compare the end-to-end training efficiency of Memo and two baselines. Table 3 shows the MFU and TGS of DeepSpeed-Ulysses, Megatron-LM and Memo under different training workloads. During evaluation, we manually adjust the distributed parallelism strategies for each system and each workload to achieve optimal training performance for fair comparisons. Detailed parallelism strategies are provided in Appendix A.
@@
Overall, Memo is capable of training longer sequences than the competitors. Across the training of 7B, 13B, 30B, and 65B models on 8, 16, 32, and 64 GPUs, DeepSpeed-Ulysses supports sequence lengths of 256K, 256K, 128K, and 1280K, while Megatron-LM supports sequence lengths of 640K, 640K, 384K, 512K. In comparison, Memo achieves superior performance in all scenarios, enabling training sequence lengths of 1024K, 1408K, 1280K, and 1408K. Megatron-LM only supports up to 640K sequence length, even if we have leveraged a high model parallel degree and enabled the memory reduction techniques. This is unsurprising since it overlooks the memory fragmentation issue, leading to OOM for large sequence lengths. DeepSpeed, thanks to its support of DeepSpeed-Ulysses sequence parallel and ZeRO-3 optimizer, is capable of training 1280K sequence length when training the 65B model on 64 GPUs. When training smaller models, DeepSpeed supports only very small sequence lengths. This is because it can only utilize a small SP size of 8, which either aligns to the number of GPUs or is dividable by the number of attention heads (40 and 56). In contrast, by token-wise recomputation/swapping and memory planning, Memo is able to train sequences over 1 million tokens in all scenarios.
@@
Furthermore, when comparing Memo to the baselines with aligned sequence lengths, Memo achieves superior MFU and TGS. Across all experimented workloads, Memo achieves an average MFU of 51.33%. In contrast, Megatron-LM and DeepSpeed only achieve an average MFU of 23.91% and 23.26%, respectively. In average, Memo achieves $2.42\times$ and $2.26\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
The deficiencies of baselines are not surprising — due to their unsatisfactory memory management, Megatron-LM and DeepSpeed require high model parallel size (i.e. large TP size and/or large SP size) and full activation recomputation to avoid OOM error, which leads to low training efficiency. For example, in order to train the 65B model with 1024K sequence length on 64 GPUs, DeepSpeed has to simultaneously use an SP size of 64, ZeRO-3 and activation recomputation to avoid OOM error. The large SP/TP size and ZeRO degree introduce significant communication overheads, while vanilla memory reduction techniques incurs extra overhead, further diminishing training efficiency. Memory fragmentation is another factor leading to the low training efficiency of these systems. When the GPU memory becomes highly fragmented and cannot allocate sufficient space for new tensors, the allocator will call expensive “cudaMalloc” and “cudaFree” operations to reorganize the GPU memory, which blocks GPU computation substantially. When training the 7B model on 8 GPUs using Megatron-LM, the memory reorganization operation is triggered 6 times and 16 times per iteration for sequence lengths of 128K and 256K, respectively.
In contrast, Memo successfully addresses these issues. On the one hand, the fine-grained activation recomputation and swapping technique significantly reduces computation burden compared to the vanilla full recomputation. This advantage is even more pronounced during long-context training. On the other hand, Memo adopts an optimized memory planning for activation, which avoids GPU memory fragmentation and the time-consuming memory reorganization process. Reducing memory fragmentation also allows for more efficient model parallelism configurations. For example, when training the 30B model on 32 GPUs with 256k sequence length, Memo adopts a TP size of 8, a CP size of 2, and a DP size of 2, which is more efficient than Megatron-LM with a TP size of 8 and CP size of 4. Additionally, Memo does not trigger any memory reorganization since the memory is already managed. As a result, Memo consistently achieves an MFU of approximately 50% across all model sizes and sequence lengths, enabling the efficient training of significantly longer sequences compared to the baselines.
@@
Memo (Fine-grained Recomputation & Swapping + Memory Plan) 47.99% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71%

### v2

@@ 4.2. Bi-level Memory Planning
The offline DSA problem handles a sequence of memory allocations and deallocations, and aims to determine the address of each allocated memory block and at the same time minimizing the peak memory usage. Parameters of offline DSA problem includes:
@@
Here the first constraint and the last two constraints define and limit peak memory, while the second and third constraints ensure non-overlapping tensors. Following this formulation, the solution for each tensor’s address is optimal. However, modern LLM training involves thousands of allocation and deallocation requests within a single training iteration, which makes this NP-hard MIP problem computationally intractable. Consequently, it’s infeasible to solve this MIP problem in one pass, given the prohibitively high time cost.
Fortunately, all transformer layers have identical structures and memory request sequences, which presents repetitive substructures within the MIP problem. By leveraging this inherent repetitiveness, we can instead devise a bi-level hierarchical MIP optimization algorithm, which is both computationally feasible and effective.
@@
As shown in the bottom of Figure 7, we first solve the offline DSA sub-problem for just one transformer layer’s forward (backward) pass, which is called the first-level MIP. The scale of the first-level MIP is small enough to tackle. This offline DSA problem can be simply solved by any MIP solver (e.g. Gurobi (Gurobi Optimization, LLC, 2024)). After this step, the peak memory needed for the forward (backward) propagation of a single transformer layer, as well as the address of each transient tensor within a transformer layer is determined. After solving the sub-problem for one transformer layer, all other transformer layers can reuse the same memory address for (de)allocation.
Besides identical transformer layers, LLMs also have other layers that process input tokens and classify output tokens. Therefore, after solving the first MIP for one-layer DSA problem, we conduct the second MIP for the whole LLM training to generate the peak memory requirement and addresses of all transient activation tensors. To simplify the optimization process, we can replace the original fine-grained memory request sequence of each transformer layer’s forward (backward) propagation with a “pseudo” large memory request pair, as shown in Figure 7. The size of this “pseudo” memory block corresponds to the memory usage of each transformer layer as determined by the first-level MIP. After the substitution, this reformulated memory request sequence also satisfies the formulation of an offline DSA problem, with a size small enough to be efficiently solved. We then leverage the MIP solver again to solve this second-level MIP problem. After this step, the addresses of all activation tensors, and the peak memory needed for all transient activation tensors can be determined.
@@ 4.3.1. Overview
Figure 9 illustrates the overall architecture of Memo. First, the job profiler takes in the model configuration, then executes a training iteration to profile the memory requests directed to the PyTorch CUDA allocator during the training phase. The job profiler also determines offloading fraction $\alpha$ by solving optimization problem in Section 4.1. These memory requests comprise a sequence of allocation and deallocation instructions. Afterwards, the memory planner receives the memory requests, executes the bi-level MIP optimization algorithm and, generates a memory plan, which constitutes the addresses of all transient activation tensors during one training iteration. Finally, the runtime executor reads the memory plan and conducts the training process.
@@ 4.3.2. Job Profiler
However, naively recording all memory requests may lead to OOM error. For example, directly profiling a GPT-7B model with a sequence length of 512K on 8 GPUs can result in OOM error. Fortunately, all transformer layers have identical memory footprint. We leverage this property by only profiling one transformer layer’s memory footprint and then applying it to all transformer layers.
When the sequence is too long, we cannot even profile one single transformer layer. In such extreme cases, we turn to the CUDA Unified Memory feature, which enables the swapping between GPU memory and CPU memory under the hook, creating an illusion of unlimited GPU memory. By integrating CUDA Unified Memory support into the PyTorch CUDA allocator, we have successfully managed to profile the training of extremely long context lengths.
The profiler also gathers the basic information to determine $\alpha$ in Section 4.1, including the size of each skeletal activation tensor, and the forward time of a one layer. Subsequently, it solves for the optimal $\alpha$ to maximize the overlapping of computation and transmission as well as to avoid CPU OOM error.
@@ 4.3.3. Memory Planner
Given the memory request sequence generated by the job profiler, memory planner executes the bi-level MIP optimization algorithm as introduced in Section 4.2 to generate a memory plan, which includes the address of each transient activation tensor and the peak memory usage needed during training. Memo uses the Gurobi (Gurobi Optimization, LLC, 2024) optimizer to solve the MIP problems. In all our experiments, memory planning takes less than 5 minutes, which is negligible compared to the training time of LLMs.
@@ 4.3.4. Runtime Executor
The runtime executor takes the memory plan, and executes the training process. It is built on the top of Megatron-LM (Shoeybi et al., 2019) with TransformerEngine (NVIDIA, 2024b), one of the most popular training frameworks. The runtime executor utilizes two rounding buffers for the storage of skeletal activations, as introduced in Section 4.1. Meanwhile, the transient activation tensors are (de)allocated according to the memory plan.
Three CUDA streams are employed for efficient overlapping of data transmission and GPU computation, which are for GPU computation, activation offloading from GPU to CPU, and activation prefetching from CPU to GPU, respectively. Figure 10 shows the scheduling of computation and transmission. After the computation of one layer’s forward pass, the skeletal activations of this layer are scheduled to be transferred to the CPU memory, which can overlap with the computation of the next layer. Before the backward computation of one layer, the forward skeletal activations of the previous layer are scheduled to be fetched back to GPU. In addition, token-wise tensor recomputation is scheduled before the layer’s backward pass. By hiding the activation swapping with computation and enabling the lightweight, token-wise activation recomputation, Memo minimizes the overhead of activation rematerialization at full stretch.
@@
5. Experiments
In this section, we conduct experiments across various model sizes and input sequence lengths to show that Memo achieves superior efficiency in long context training of LLMs.
5.1. Setup
Baselines: We select two widely-used LLM training frameworks as baselines. The first is Megatron-LM (commit id: ccfeda47cb) (Shoeybi et al., 2019) in conjunction with TransformerEngine (v1.3) (NVIDIA, 2024b). Megatron-LM, maintained by NVIDIA, is renowned for its comprehensive support of hybrid parallelisms, including DP, TP, $P$, SP, and CP. The other baseline is Megatron-Deepspeed (commit id 7eb36a11b3) paired with DeepSpeed (v0.14.3) (Rasley et al., 2020), which is recognized for ZeRO optimizers and DeepSpeed-Ulysses (Jacobs et al., 2023), a novel parallel training strategy designed for long context LLM training. For all baselines, we employ FlashAttention (Dao et al., 2022; Dao, 2024) and mixed-precision training (Micikevicius et al., 2018; Kalamkar et al., 2019), which are commonly used in LLM training.
Metrics: We use three important evaluation metrics to measure the training efficiency: Model FLOPs Utilization (MFU), Tokens per GPU per Second (TGS), and wall-clock time. MFU is defined as the ratio of model FLOPs per second to the theoretical peak FLOPs per second of the GPU (e.g. 312 TFLOPS for NVIDIA A800 GPUs) (Korthikanti et al., 2022). Based on FlashAttention (Dao et al., 2022; Dao, 2024) and considering the causal mask, the exact formula for calculating model FLOPS per sample is:
$6\cdot s\cdot P+6\cdot n\cdot h\cdot s^{2}.$
MFU is a standard metric that measures the training efficiency of how model FLOPs utilize computational resources. On the other hand, TGS directly measures training throughput, and wall-clock time reflects the latency of training each batch, providing a clear view of how quickly a model can be trained using a given amount of training samples. All metrics are crucial for LLM researchers and engineers, enabling comparisons among various training strategies (including distributed parallelisms and activation recomputation).
@@ 5.1. Setup
Workloads: Our experiments cover a wide range of workloads to examine the strength of Memo. In particular, we consider training the 7B, 13B, 30B and 65B GPT models on 8, 16, 32, and 64 GPUs, respectively, with various sequence lengths ranging from 4K to 1408K. The detailed model configurations are shown in Table 2.
@@
Table 3. MFU (Model FLOPS Utilization), TGS (Tokens per GPU per Second), and wall-clock time of different methods on different numbers of GPUs, model sizes, and sequence lengths. “DS” stands for “DeepSpeed”, and “Mega” stands for “Megatron-LM”. ✗oom stands for Out Of Memory error when the system runs out of GPU memory, and ✗oohm stands for Out Of Host Memory error, which means the CPU memory is depleted. “h”, “m” and “s” represent hour, minute and second respectively.
7B GPT on 8 GPUs 13B GPT on 16 GPUs 30B GPT on 32 GPUs 65B GPT on 64 GPUs
SeqLen DS Mega Memo DS Mega Memo DS Mega Memo DS Mega Memo
44.71% 49.45% 49.45% 44.08% 46.17% 46.17% 43.70% 50.67% 50.67% 43.89% 53.61% 53.61%
3235.64 3578.86 3578.86 1675.53 1755.23 1755.23 724.44 840.14 840.14 337.95 412.90 412.90
4K 2.53s 2.29s 2.29s 2.44s 2.33s 2.33s 2.83s 2.44s 2.44s 3.03s 2.48s 2.48s
46.28% 51.57% 51.57% 38.90% 50.91% 50.91% 38.78% 49.27% 49.27% 43.34% 53.27% 53.27%
3116.25 3472.95 3472.95 1393.24 1823.52 1823.52 615.22 781.68 781.68 321.00 394.61 394.61
8K 5.26s 4.72s 4.72s 5.88s 4.50s 4.50s 6.66s 5.24s 5.24s 6.38s 5.19s 5.19s
36.41% 51.29% 51.29% 36.95% 51.52% 51.52% 34.54% 49.84% 49.84% 40.98% 51.10% 51.10%
2152.65 3032.14 3032.14 1186.24 1654.11 1654.11 504.64 728.22 728.22 282.05 351.58 351.58
16K 15.2s 10.8s 10.8s 13.8s 9.91s 9.91s 16.2s 11.2s 11.2s 14.5s 11.6s 11.6s
35.91% 52.75% 52.75% 34.51% 49.79% 49.79% 35.19% 48.06% 48.06% 35.73% 34.00% 44.79%
1706.13 2506.39 2506.39 917.96 1324.27 1324.27 443.77 606.16 606.16 215.32 204.83 269.84
32K 38.4s 26.1s 26.1s 35.7s 24.7s 24.7s 36.9s 27.0s 27.0s 38.0s 39.9s 30.4s
27.95% 41.55% 52.34% 27.97% 38.51% 52.65% 29.93% 35.76% 52.12% 31.05% 22.79% 47.80%
935.76 1417.74 1786.22 553.85 762.46 1042.50 296.41 354.20 516.16 149.80 109.94 230.62
64K 2m17s 1m33s 1m13s 1m58s 1m26s 1m3s 1m51s 1m33s 1m3s 1m49s 2m19s 1m11s
25.46% 24.13% 50.96% 25.45% 23.02% 50.93% 25.54% 14.70% 49.66% 26.13% 15.10% 48.61%
555.51 526.62 1111.99 333.51 301.67 667.41 176.92 101.85 344.09 90.15 52.10 167.69
128K 7m52s 8m18s 3m56s 6m33s 7m15s 3m16s 6m10s 10m43s 3m10s 6m4s 10m30s 3m15s
23.38% 29.07% 53.62% 21.98% 25.30% 51.22% oom 17.15% 50.00% 22.07% 9.57% 49.87%
296.48 368.65 679.92 171.82 197.78 400.39 oom 72.24 216.41 48.51 21.03 109.58
256K 29m28s 23m42s 12m51s 25m26s 22m5s 10m55s oom 29m30s 10m6s 22m31s 51m56s 9m58s
oom 27.98% 53.04% oom 22.88% 51.91% oom 23.32% 50.69% 20.40% 12.07% 48.85%
oom 250.07 474.02 oom 127.40 289.11 oom 73.37 159.52 32.89 19.46 78.76
384K oom 52m25s 27m39s oom 51m26s 22m40s oom 44m39s 20m32s 49m49s 1h24m 20m48s
oom 34.43% 51.84% oom 29.10% 52.40% oom oom 51.06% 19.83% 5.32% 49.71%
oom 237.56 357.70 oom 125.86 226.65 oom oom 126.23 25.24 6.77 63.29
512K oom 1h14m 48m51s oom 1h10m 38m33s oom oom 34m37s 1h27m 5h23m 34m31s
oom 30.90% 52.59% oom 19.41% 52.13% oom oom 51.72% 19.06% oom 50.05%
oom 173.63 295.51 oom 83.94 184.33 oom oom 105.28 20.04 oom 52.65
640K oom 2h6m 1h14m oom 2h10s 59m15s oom oom 51m52s 2h16m oom 51m52s
oom oom 51.89% oom oom 51.71% oom oom 51.18% 19.53% oom 51.16%
oom oom 245.76 oom oom 154.63 oom oom 88.55 17.50 oom 45.84
768K oom oom 1h47m oom oom 1h25m oom oom 1h14m 3h7m oom 1h11m
oom oom 52.71% oom oom 51.76% oom oom 51.50% 19.12% oom 51.05%
oom oom 251.98 oom oom 134.08 oom oom 77.48 14.93 oom 39.85
896K oom oom 2h22m oom oom 1h54m oom oom 1h39m 4h16m oom 1h36m
oom oom 52.30% oom oom 52.06% oom oom 51.24% 19.00% oom 51.27%
oom oom 188.73 oom oom 118.96 oom oom 68.20 13.14 oom 35.45
1024K oom oom 3h5m oom oom 2h27m oom oom 2h8m 5h33m oom 2h3m
oom oom oom oom oom 51.74% oom oom 51.73% 19.11% oom 51.20%
oom oom oom oom oom 105.74 oom oom 61.72 11.86 oom 31.77
1152K oom oom oom oom oom 3h6m oom oom 2h39m 6h43m oom 2h35m
oom oom oom oom oom 51.78% oom oom 51.59% 18.90% oom 51.42%
oom oom oom oom oom 95.72 oom oom 55.79 10.64 oom 28.94
1280K oom oom oom oom oom 3h48m oom oom 3h16m 8h33m oom 3h9m
oom oom oom oom oom 52.10% oom oom oohm oom oom 51.45%
oom oom oom oom oom 87.93 oom oom oohm oom oom 26.50
1408K oom oom oom oom oom 4h33m oom oom oohm oom oom 3h47m
@@ 5.2. End-to-end Evaluation
We compare the end-to-end training efficiency of Memo and two baselines. Table 3 shows the MFU, TGS, and wall-clock time of DeepSpeed-Ulysses, Megatron-LM and Memo under different training workloads. During evaluation, we manually adjust the distributed parallelism strategies for each system and each workload to achieve optimal training performance for fair comparisons.
@@
Overall, Memo is capable of training longer sequences than the competitors. Across the training of 7B, 13B, 30B, and 65B models on 8, 16, 32, and 64 GPUs, DeepSpeed-Ulysses supports sequence lengths of 256K, 256K, 128K, and 1280K, while Megatron-LM supports sequence lengths of 640K, 640K, 384K, 512K. In comparison, Memo achieves superior performance in all scenarios, enabling training sequence lengths of 1024K, 1408K, 1280K, and 1408K. Megatron-LM only supports up to 640K sequence length, even if we have leveraged a high model parallel degree and enabled the memory reduction techniques. This is unsurprising since it overlooks the memory fragmentation issue, leading to OOM for large sequence lengths. DeepSpeed, thanks to its support of DeepSpeed-Ulysses sequence parallel and ZeRO-3 optimizer, is capable of training 1280K sequence length when training the 65B model on 64 GPUs. When training smaller models, DeepSpeed supports only very small sequence lengths. This is because it can only utilize a small SP degree of 8, which either aligns to the number of GPUs or is dividable by the number of attention heads (40 and 56). In contrast, by token-wise recomputation/swapping and memory planning, Memo is able to train sequences over 1 million tokens in all scenarios.
When training with relatively short sequences ranging from 4K to 32K, there’s little memory pressure. In such cases, activation recomputation is often unnecessary, and different methods may use the same parallelism strategies. Consequently, Memo falls back to Megatron-LM, resulting in the same performance. Nevertheless, Memo significantly outperforms the baselines when the sequence length reaches 64K or even higher.
@@
Furthermore, when comparing Memo to the baselines with aligned sequence lengths, Memo achieves superior MFU, TGS, and wall-clock time. Across all experimented workloads, Memo achieves an average MFU of 51.04%. In contrast, Megatron-LM and DeepSpeed only achieve an average MFU of 35.01% and 30.74%, respectively. On average, Memo achieves $1.97\times$ and $1.80\times$ MFU compared to Megatron-LM and DeepSpeed, respectively.
The deficiencies of baselines are not surprising — due to the unsatisfactory memory management, they are forced to utilize less-efficient configurations of parallelism strategies and rematerialization options. [^6] We have provided the detailed configurations in our supplementary materials. Firstly, the baselines need to employ a high model parallelism degree (i.e. large TP degree and/or large SP degree) to avoid OOM error. For instance, when training the 65B model with 256K sequence length, Megatron-LM has to use a TP degree of 16, while Memo can support a TP degree of 8. Given the much lower inter-node bandwidth compared to intra-node, the TP communication cost of Memo is 63% lower than that of Megatron-LM (44.3 v.s. 120.4 seconds). Secondly, the baselines require full activation recomputation to avoid the OOM error, leading to over 20% redundant GPU calculation. In contrast, Memo integrates fine-grained recomputation and swapping, incurring smaller overhead. Last but not least, the lack of memory planning often triggers memory reorganization during training, which blocks GPU computation and significantly hampers training efficiency. When training the 7B model on 8 GPUs using Megatron-LM, the memory reorganization operation is triggered 6 times and 16 times per iteration for sequence lengths of 128K and 256K, respectively. A detailed ablation study is presented in Section 5.3.
As a result, Memo consistently achieves an MFU of approximately 50% across all model sizes and sequence lengths, enabling more efficient training of significantly long sequences compared to the baselines.
@@
Memo (Fine-grained Management + Memory Plan) 47.99% 50.96% 53.62% 53.04% 51.84% 52.59% 51.89% 52.71%
Table 5. Impact of offloading fraction $\alpha$ on training efficiency.
Value of $\alpha$
Sequence Length 0.000 0.125 0.250 0.375 0.500 0.625 0.750 0.875 1.000
192K 51.04% 51.86% 52.14% 52.30% 52.72% 52.78% 53.11% 51.74% 49.69%
256K 52.03% 52.05% 52.19% 52.45% 52.70% 52.96% 53.21% 53.29% 53.62%
320K 52.15% 52.36% 52.52% 52.66% 52.92% 53.10% 53.35% oohm oohm
384K 52.09% 52.31% 52.45% 52.58% 53.04% oohm oohm oohm oohm
5.3. Ablation Studies
Next, we assess the effectiveness of the proposed techniques in Memo. All experiments in ablation studies are conducted by training the 7B model on 8 GPUs, keeping the parallelism configuration fixed at a TP degree of 4 and a CP degree of 2.
5.3.1. Effectiveness of Memory Planning
Additionally, the frequent GPU memory reorganization process further impairs training efficiency. By employing memory planning, the fragmentation issue can be eliminated, providing more memory for longer context training. Getting rid of GPU memory reorganization, memory planning brings an average of $1.51\times$ MFU when facing the same context length.
5.3.2. Effectiveness of Token-wise Recomputation
For token-wise recomputation and swapping, we compare Memo and its variants, one with full recomputation and another with full swapping. The results are shown in the last three rows of Table 4. When training with appropriate sequence length, which is 256k in this scenario, the computation time of one transformer layer can fully overlap with the offloading time of a layer’s activations. Therefore, full swapping with memory planning can achieve an MFU of 53.62% under 256K sequence length, far exceeding the 42.05% achieved by full recomputation with memory planning. However, for short sequence lengths, such as 64K, the offloading time of one layer’s activations blocks the GPU computation, resulting in a lower MFU than full recomputation. Full swapping presents another challenge as the sequence length grows longer: the host memory is rapidly depleted by offloaded activations, leading to OOHM errors. Full recomputation, which does not offload the input tensor of each layer, also encounters OOM errors at a sequence length of 768K.
By employing token-wise recomputation together with swapping, Memo consistently improves training efficiency for both short and long context lengths. For short sequence lengths like 64K, our tensor-level design only offloads the input tensor of the transformer layer and the FlashAttention output tensor to CPU memory, enabling efficient overlap of GPU computation and data transmission. For long context lengths, our rounding-buffer token-level management successfully avoids depleting the CPU memory, and incurs only minimal recomputation overhead. Among all the methods, Memo supports the longest sequence length. Considering MFU, Memo achieves an average of $1.22\times$ MFU compared to full recomptutation with memory planning, and an average of $1.13\times$ MFU compared to full offloading with memory planning.
5.3.3. Effect of the Offloading Fraction $\alpha$
To investigate the impact of the offloading fraction $\alpha$ on training efficiency, we train a 7B model on 8 GPUs, with TP=4=4, CP=2=2 and varying $\alpha$ values. The results are presented in Table 5. For a sequence length of 192K, the MFU initially increases with $\alpha$, peaking at 53.11% when $\alpha=0.75$, and then declines. This trend occurs because, for $\alpha<0.75$, a higher $\alpha$ reduces recomputation overhead by increasing the proportion of skeletal activations that are swapped. Beyond this point, however, further increases in $\alpha$ lead to swapping overhead that exceeds the computation time of an individual transformer layer, stalling GPU computations until the swapping concludes and thus reducing the MFU. At a sequence length of 256K, the computation time of a single transformer layer, due to the quadratic complexity of self-attention, surpasses the offloading time for all skeletal activations in this layer. Consequently, training efficiency consistently improves as $\alpha$ increases. For longer sequences such as 320K and 384K, offloading all skeletal activations becomes unviable due to CPU memory constraints. Considering the constraints, Memo can effectively identify the optimal $\alpha$ to minimize recomputation overhead.